the height of a triangle is 2 more than twice the base. the area is 20 square feet. find the base.
Answer by stanbon(57277)

the height of a triangle is 2 more than twice the base. the area is 20 square feet. find the base.
--------------
Area = (1/2)(base)(height)
-------------------------
20 = 1/2 (x)(2x+2)
40 = 2x^2+2x
20 = x^2+x
x^2+x-20 = 0
(x-5)(x+4) = 0
x = 5 or x=-4
----------------------
Only x=5 ft. is a possible answer for the base:
Then height = 2*5+2 = 12 ft.
========================
Cheers,
Stan H.